OpenStudy (anonymous):

check my answer Which equation best represents the following scenario? Heidi answered 4 more questions correctly on the quiz than Sean. Heidi answered 8 questions correctly. four s equals eight s plus four equals eight four minus s equals eight eight divided by s equals four i got the third one explain how u got your answer and then i will medal and fan you

OpenStudy (anonymous):

its the second one

OpenStudy (anonymous):

@jameswolf how did you get the second one

OpenStudy (anonymous):

so s is the number sean got, h is the number heidi got h = 8 h - 4 = s replace 8 from the first, into the second 8 - 4 = s add 4 to both sides 8 = s + 4

OpenStudy (anonymous):

thank you @Jameswolf